• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 07:58:42 on April 6,2011

    Entry number 354567

    keyword=Owl shift summary

    Shift started with HRS and DAQ problems and a controlled access requested.

    Problems with HRS were solved by 1am. IOC22 was rebooted. Also, ROC5 was physically powercycled in the Hall. We were back in production by 1:30AM.

    1:00 Hall C inserted in halfwave plate (it is now IN).

    5:30 MCC called notifying a problem with cryo epics. Beam down.

    6:30 Beam is back. Production resumed.

    Shift finished running production (run 3317)


    Runlist
    ========
    Run BCM_U1(C) SUM(C)
    3312 0.00196 0.00196 4He
    3313 0.01802 0.01998 4He
    3314 0.01811 0.03809 4He
    3315 0.01706 0.05515 4He
    3316 0.01798 0.07313 4He
    Sum BCM U1 (C) = 0.07313